An electric utility is considering a new power plant in northern Arizona. Power from the plant would be sold in the Phoenix area, where it is badly needed. Because the firm has received a permit, the plant would be legal; but it would cause some air pollution. The company could spend an additional $40 million at Year 0 to mitigate the environmental problem, but it would not be required to do so. The plant without mitigation would cost $240.62 million, and the expected cash inflows would be $80 million per year for 5 years. If the firm does invest in mitigation, the annual inflows would be $84.19 million. Unemployment in the area where the plant would be built is high, and the plant would provide about 350 good jobs. The risk-adjusted WACC is 16%.

Calculate the NPV and IRR with mitigation. Enter your answer for NPV in millions. For example, an answer of $10,550,000 should be entered as 10.55. Negative values, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to two decimal places.

NPV: $ million

IRR: %

Calculate the NPV and IRR without mitigation. Enter your answer for NPV in millions. For example, an answer of $10,550,000 should be entered as 10.55. Negative values, if any, should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to two decimal places.

NPV: $ million

IRR: %
Answer #1

NPV and IRR with mitigation:

NPV = -$4.96 million and IRR = 15.24%

NPV computation:

Year Cash flow 1+r PVIF PV = PVIF*cash flow
0 -280.62 1.16 1.0000 -280.62
1 84.19 0.8621 72.58
2 84.19 0.7432 62.57
3 84.19 0.6407 53.94
4 84.19 0.5523 46.50
5 84.19 0.4761 40.08
NPV -4.96

IRR computation:

Year Cash flow 1+r PVIF PV = PVIF*cash flow
0 -280.62 1.1524 1.0000 -280.62
1 84.19 0.8678 73.06
2 84.19 0.7530 63.39
3 84.19 0.6534 55.01
4 84.19 0.5670 47.74
5 84.19 0.4920 41.42
NPV 0.00000

Without mitigation:

NPV = $21.32 million and IRR = 19.74%

NPV calculations:

Year Cash flow 1+r PVIF PV = PVIF*cash flow
0 -240.62 1.16 1.0000 -240.62
1 80.00 0.8621 68.97
2 80.00 0.7432 59.45
3 80.00 0.6407 51.25
4 80.00 0.5523 44.18
5 80.00 0.4761 38.09
NPV 21.32

IRR computation:

Year Cash flow 1+r PVIF PV = PVIF*cash flow
0 -240.62 1.1974 1.0000 -240.62
1 80.00 0.8351 66.81
2 80.00 0.6975 55.80
3 80.00 0.5825 46.60
4 80.00 0.4864 38.92
5 80.00 0.4062 32.50
NPV 0.00000